OpenStudy (anonymous):

\[r=2\sqrt{5L}\]the formula can be used to approximate the speed R in miles per hour of a car that has left skid marks of length L in feet how far will a car skid at 70mph?

OpenStudy (anonymous):

L=245 feet at 70 mph

OpenStudy (stormfire1):

\[70/2 = \sqrt{5L}\]\[35=\sqrt{5}\sqrt{L}\]\[35^{2}=5*L\]\[1225/5=L\]L=245ft.